[figure: 9.7 cm]

LOW TEMPERATURE TRANSPORT AND MAGNETIC PROPERTIES OF INTERMEDIATE VALENT SmB_{6}

S. Gabáni^{a}, K. Flachbart^{b}, V. Pavlík^{b}, E. Konovalova^{c}, Y. Paderno^{c} and

T. Herrmannsdörfer^{d}

aDepartment of Experimental Physics, Šafárik University, 04154 Košice, Slovakia

bInstitute of Experimental Physics, SAS, 04353 Košice, Slovakia cInstitute for Problems of Materials Science, UAS, 252680 Kiev, Ukraine dPhysikalisches Institut, Universität Bayreuth, 95440 Bayreuth, Germany

We present results of transport and magnetic properties of three single-crystalline samples of the intermediate valent small-gap semiconductor SmB_{6} at low temperatures. The received resistivity dependences of all samples below 0.5 K exhibit an activated behavior with an energy gap of a few mK. The temperature dependences of magnetic susceptibility show an increase below 15 K which can be accounted for by impurities, bare Sm^{3+} ions or by a small amount of magnetic 4f^{5}5d^{1} in-gap configurations of Sm-ions.

[figure: 13.4 cm]

Subject category :

1. Correlated Electrons and High Temperature Superconductors

Presentation mode : poster

Corresponding author : S. Gabáni

Address for correspondence : Mgr. Slavomír Gabáni Department of Experimental Physics Faculty of Sciences P. J. Šafárik University Park Angelinum 9 SK-04154 Košice, Slovakia phone: +421 95 6221128 fax: +421 95 6336292

Email address : gabani@kosice.upjs.sk